The problem describes a debt to be amortized. (Round your answers to the nearest cent.) A man buys a house for $350,000. He makes a $150,000 down payment and amortizes the rest of the purchase price with semiannual payments over the next 9 years. The interest rate on the debt is 12%, compounded semiannually.

(a) Find the size of each payment. (b) Find the total amount paid for the purchase. (c) Find the total interest paid over the life of the loan.
Answer #1

First we find the amount of installment to be paid semiannually using the financial calculator

Feed N = 9 * 2 = 18 period

IY = 6 for each period

PV = 350000 - 150000 = 200000

FV = 0

Compute PMT we get 18471.31

a) Size of payment = 18471.31

b) Total Payment = 18 * 18471.31 = 332,483.55

c) Total Interest Payment = Total Payment - Original Loan Amount

= 332483.55 - 200000

= 132483.55